OpenStudy (anonymous):

Given that F(x) = x 2 + 2, evaluate F(1) + F(5). @phi @popcornfun

OpenStudy (phi):

do you mean F(x) = x^2 + 2 i.e \[ F(x) =x^2 +2 \]

OpenStudy (anonymous):

these are the answer choices @phi 38 30 15

OpenStudy (phi):

do you mean F(x) = x^2 + 2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@ph

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@MDoodler

OpenStudy (anonymous):

u still need help?

OpenStudy (anonymous):

@tmoneycash3